OpenStudy (anonymous):

Someone please help me out!! I need to figure this out as soon as possible! I don't get geometry! :( please? What is the slope-intercept form of the equation y-7= -5/2(x+4)?

OpenStudy (anonymous):

im good at geometry

OpenStudy (anonymous):

I'm glad. Please help me out????

OpenStudy (anonymous):

haha okay whats the answer they give u

OpenStudy (anonymous):

There are four of them and in this one, my teacher circled the answer but i have to explain to him how to get to the answer and that's why i need to show work. I didn't get it right the first time. 1.) y-2=-5/2(x+2) 2.) y+7=-x+5/2 3.) y=-4/7+2 4.) y=-5/2x-3 my teacher says that the 4th one is correct, but how to i get the answer????? that's the confusing part.

OpenStudy (anonymous):

okay wait i need to see something real quick

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i working it out on a paper so it might be a while sorry and what is the slop intercept form

OpenStudy (anonymous):

slope intercept form is y=mx+b.

OpenStudy (anonymous):

okay first you need to write this in slope intercept for y-7=-5/2 (x+4)

OpenStudy (anonymous):

okay. got that. is this is answer? I'm just really desperate. Today is teh last day before the semester ends so i wanna redo this test to get my grade up.

OpenStudy (anonymous):

this is not answer im working on it

OpenStudy (anonymous):

okay. sorry

OpenStudy (anonymous):

then what you would do next is use the distributive property

OpenStudy (anonymous):

which mean u multiply everything in the parenthesis by -5/2

OpenStudy (anonymous):

you would get 5/2x and then -5/2 times 4 is 20/2

OpenStudy (anonymous):

then you take the 20/2 and divide it and you get 10

OpenStudy (anonymous):

okay. got that so far.

OpenStudy (anonymous):

right now ur equation should look like y-7=-5/2x-10

OpenStudy (anonymous):

after tht u would subtract 7 from each side so ur answer would be y=-5/2x-3

OpenStudy (anonymous):

omg! you're a mircale worker! Thanks so much!!! :) I'm starting to get it now!! God bless you 1000 times!

OpenStudy (anonymous):

haha ur welcome